# It’s Giving Major Satisfaction: Why Fairness Matters for Developers

### v1

@@ V-C Grouping Job Satisfaction Factors
To uncover these natural groupings, we first checked for the independence of factors by calculating the correlations between the satisfaction scores for each pair of factors using the Spearman method. The Spearman correlation method is suitable for ordinal data, like Likert-scale responses, because it assesses how well the relationship between two variables can be described by a monotonic function [75]. For example, if the satisfaction scores for feedback and appreciation are highly correlated, it suggests that respondents who rate feedback highly also tend to rate appreciation highly.
After establishing the relationships between factors through correlation analysis, we computed the dissimilarity matrix using the Manhattan distance, which is appropriate for ordinal data. The Manhattan distance measures the absolute differences between the pairwise correlation coefficients, summing these differences across all pairs [76]. This method helps to quantify how dissimilar each pair of job satisfaction factors is based on their correlation values [76]. For instance, if the Manhattan distance between satisfaction with benefits and satisfaction with salary is low, it suggests these factors are closely related in how respondents perceive them.
We then performed hierarchical clustering using the average linkage method, implemented in R with the hclust() function [77, 78]. The average linkage method calculates the average distance between all pairs of factors in two clusters, merging the clusters with the smallest average distance first. The result of the hierarchical clustering is visualized in a dendrogram (see Figure 1), a tree-like diagram that illustrates how clusters are merged at different distance thresholds. For instance, if satisfaction with benefits and satisfaction with salary are closely related, they will be merged early in the dendrogram.
@@
Additionally, we made a slight adjustment by separating the overall satisfaction factor from CF3, treating it as a distinct composite factor. This adjustment led to a final set of seven composite factors, as shown in Table III. For example, satisfaction with benefits and salary was highly correlated, so we combined these factors into the composite factor compensation. When combining factors, we averaged their scores; for instance, if the satisfaction scores for benefits and salary were $3$ and $4$, respectively, the score for compensation would be $(3+4)/2=3.5$.
@@
Figure 1: Dendogram of clustering job satisfaction factors into groups.
Table III: Composite Factors of Practitioners’ Job Satisfaction.
@@
V-D Ordinal Logistic Regression Analysis
After establishing the job satisfaction clusters, we aimed to understand how fairness perception or fairness contexts are associated with the probability of job satisfaction being in a higher or lower category. To achieve this, we performed an ordinal logistic regression analysis. Ordinal logistic regression predicts the probability of an observation falling into one of several ordered categories, making it ideal for analyzing outcomes like job satisfaction levels [79].
@@ V-D Ordinal Logistic Regression Analysis
V-E Moderation Analysis
After performing the ordinal logistic regression analysis, we explored how the observed relationships differ across demographic groups by employing moderation analysis through RQ2 and RQ3.1. Moderation analysis tests whether a variable, such as a demographic factor, influences the direction and/or strength of the relationship between an independent variable (e.g., perception of fairness) and a dependent variable (e.g., job satisfaction) [80]. For example, suppose we find that perceptions of procedural fairness are positively associated with job satisfaction. A moderation analysis might examine whether this relationship is stronger for women than for men.
@@ V-E Moderation Analysis
In our study, we specifically examined the potential moderating effects of various demographic factors, as described in Section VI-A. Due to the imbalance in the distribution of demographic categories among our respondents (as shown in Table IV), initially, we were only able to effectively analyze the moderating effects for four demographic groups: gender, ethnicity representation, and work limitation. Subsequently, we simplified the categories for the remaining demographic groups to improve the analysis. For instance, ethnicity was grouped into Asian and Non-Asian categories, and work experience was divided into less than 3 years and more than 3 years. This refinement allowed us to effectively analyze the moderating effects for one additional demographic group: work experience.
However, the limited number of respondents in other remaining demographic categories, namely age, ethnicity, role and role nature restricted our ability to conduct reliable moderation analyses for these groups. This limitation is due to the potential for biased estimates caused by imbalanced distribution within the sample [81].
VI Findings
VI-A Respondents Demographics
The demographic data is summarized in Table IV. There were 80 male respondents and 25 female respondents. The age distribution shows that 100 respondents were under 40 years old, while 7 were 40 years or older. Regarding ethnicity, the majority of respondents identified as Asian (79) and White (18). In terms of job roles, most respondents were Software Developers/Engineers (72) and Software Testers/QA Engineers (12). For work experience, 11 respondents had $\leq$ 1 year of work experience, 40 had $1-3$ years, 26 had $3-5$ years, 14 had $5-7$ years, and 16 had over 7 years of experience. Regarding work limitations, 55 respondents reported no limitations, 43 indicated some form of limitation.
Table IV: Respondents’ Demographics
@@ VI-A Respondents Demographics
VI-B Survey Responses Summary
@@
VI-C RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
The results of the analysis are summarized in the first section of Table VI. We considered relationships to be impactful if they were statistically significant ($p<0.05$) and had OR exceeding 50%, to ensure that the identified factors not only demonstrate a reliable statistical relationship but also have a practically meaningful impact on job satisfaction (see [62]). Our analysis demonstrates that perceptions of fairness across various dimensions significantly shape job satisfaction among software practitioners.
@@ VI-C RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
Table VI: Ordinal Regression Analysis Result
@@
Takeaway 1: Fairness across all four dimensions; distributive, procedural, interpersonal, and informational, significantly enhances overall job satisfaction and job security for software practitioners.
VI-D RQ2: Does the relationship between fairness perceptions and job satisfaction change based on demographic groups? If so, how?
In this RQ, we aim to determine whether a software practitioner’s demographic background influences how their perceptions of fairness affect job satisfaction using moderation analysis. The demographic groups analyzed include gender, ethnicity representation, work experience and work limitation. Similar to the previous analysis, we considered a demographic group to be influential if the effect was statistically significant ($p<0.05$) and the effect size exceeded 50% (see [62]).
Our analysis reveals that the relationship between fairness perceptions and job satisfaction does indeed vary across demographic groups. Table VII (first section) presents the key findings, with different emojis indicating how the perception of fairness influences job satisfaction among practitioners in the demographic groups represented by those emojis.
@@
Distributive, procedural and informational fairness were found to have the significant impact on female developers’ satisfaction with their manager (MNG). All distributive fairness items (D1-D4) have an average 62% stronger effect on female practitioners compared to male practitioners. Additionally, other items related to procedural fairness, such as accurate information (P3), free of bias processes (P4), consistency in applying processes (P6), and appealing decisions (P7), as well as tailored communication (Inf4) from informational fairness, also contribute to female developers’ satisfaction with their manager. This suggests that to increase female developers’ satisfaction with their managers, organizations should focus on ensuring fair outcomes, accurate information, unbiased processes, and tailored communication.
It was also found that female developers’ satisfaction with job security (SEC) is influenced by all four dimensions of fairness perceptions. Key factors include receiving outcomes that accurately reflect their work (D2) and effort (D4), the ability to voice opinions (P2), access to accurate information (P3), free of bias processes (P4), the ability to appeal decisions (P7), being treated with respect (Int2) and dignity (Int3), and receiving comprehensive explanations (Inf5). To improve job security satisfaction for female developers, organizations should ensure that performance outcomes are fair and reflective of their efforts, foster a respectful environment, and maintain clear, unbiased communication.
@@
Takeaway 2: The impact of fairness perceptions on job satisfaction is stronger for female, underrepresented, less experienced practitioners, and those with work limitations. This suggests that different demographic groups have distinct needs.
@@
VI-E RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
The results of the regression analysis are summarized in the third section of Table VI. We found positive and significant correlations between job satisfaction and fairness experiences in specific SE contexts (see [62]).
@@ VI-E RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
Our analysis confirms that fairness perceptions influencing job satisfaction do indeed differ across demographic groups within SE contexts (see [62]). Table VII (last section) illustrates that fairness perceptions impacting job satisfaction do vary across demographic groups in SE contexts. Different emojis in the table indicate how perceptions of fairness influence job satisfaction among practitioners within the demographic groups represented by those emojis.
@@
This suggests that female developers feel more secure in their jobs when organizations ensure that performance outcomes are fair and reflective of their contributions [82], foster a respectful work environment [83], and maintain clear and unbiased communication [84]. These can be applied for example, when a female developer is assigned to work (i.e., allocation of work) on a critical project with tight deadlines (i.e., demand), and that workload distribution is monitored (i.e., policy) [85].
@@
We also found that unbiased processes (P4) is particularly important for underrepresented practitioners’ overall satisfaction (ALL), especially in the context of policy (C4). This suggests that ensuring unbiased processes likely helps underrepresented developers feel that they are being treated fairly, which can mitigate feelings of marginalization or exclusion [86].
@@
This suggests that how working hours are managed plays a significant role in shaping the performance satisfaction of less experienced practitioners. It could reflect how an organization or manager considers and respects a practitioner’s time and personal boundaries in relation to their workload. This includes balancing work and personal life, ensuring reasonable working hours, and handling periods of overtime with mutual respect. Less experienced practitioners may be more sensitive to these factors, as they are still establishing themselves professionally [87, 88], making respectful management of their time essential to their perception of fairness and performance satisfaction.
@@
This could be because, at this early stage in their careers, less experienced practitioners are seeking affirmation that their contributions are valued, and they rely more heavily on external validation, such as feedback, promotions, or compensation, to gauge their progress [89, 90]. In contexts such as policy, working hours, allocation of work, and demand, these practitioners are still learning how the workplace functions and what is expected of them. When their efforts lead to clear, positive outcomes, they are more likely to feel valued and motivated. On the other hand, if their contributions are not recognized or they feel disrespected, it may lead to dissatisfaction, as they might interpret this as a lack of growth.
@@
Takeaway 4: Regarding job satisfaction, female developers are more impacted by fairness in policy implementation and evaluations. Underrepresented practitioners value fairness in policy implementation and during high-demand situations. Less experienced practitioners place greater emphasis on fairness when navigating working hours, while practitioners with work limitations prioritize fairness during demanding situations.
VII Discussion
VII-A On Overall Satisfaction and Job Security
This study found that all four dimensions of fairness, distributive, procedural, interpersonal, and informational fairness, significantly influence overall job satisfaction and satisfaction with job security. Job security in this context relates to the perceived stability of one’s employment.
In the field of SE, prior research has emphasized the importance of various factors such as autonomy, task variety, and feedback in influencing job satisfaction, particularly in agile environments [91]. Autonomy, the degree to which software engineers can control how they work, and feedback on performance are particularly important in shaping their sense of satisfaction and productivity. Similarly, Storey et al. highlighted appreciation, rewards, impactful work, and work culture as important factors contributing to overall job satisfaction for developers [12].
The closest concept to job security in SE literature is turnover. Turnover in software development refers to the resignation or replacement of software developers, which may lead to severe problems like project postponement and failure [92]. In relation to turnover, studies have found that factors such as satisfaction with supervision, workplace morale, and managerial support significantly impact turnover intentions. Research on the impact of developer turnover in open-source software projects highlights the negative impact of high developer turnover, especially among external newcomers, on software quality [93].
Agile methods have also been linked to higher job satisfaction and lower turnover rates. A study showed that team members using agile methods reported greater job satisfaction and lower turnover intentions compared to those on non-agile teams [33]. Finally, job satisfaction, onboarding success, and workplace relationships are key factors influencing turnover intentions among software professionals. Training and support for new hires are crucial for retention in these contexts [94].

### v2

@@ 5.3. Grouping Job Satisfaction Factors
To uncover these natural groupings, we first checked for the independence of factors by calculating the correlations between the satisfaction scores for each pair of factors using the Spearman method. The Spearman correlation method is suitable for ordinal data, like Likert-scale responses, because it assesses how well the relationship between two variables can be described by a monotonic function (Winter et al., 2016). For example, if the satisfaction scores for feedback and appreciation are highly correlated, it suggests that respondents who rate feedback highly also tend to rate appreciation highly.
After establishing the relationships between factors through correlation analysis, we computed the dissimilarity matrix using the Manhattan distance, which is appropriate for ordinal data. The Manhattan distance measures the absolute differences between the pairwise correlation coefficients, summing these differences across all pairs (Weiß, 2019). This method helps to quantify how dissimilar each pair of job satisfaction factors is based on their correlation values (Weiß, 2019). For instance, if the Manhattan distance between satisfaction with benefits and satisfaction with salary is low, it suggests these factors are closely related in how respondents perceive them.
We then performed hierarchical clustering using the average linkage method, implemented in R with the hclust() function (Langfelder and Horvath, 2012; Guénard and Legendre, 2022). The average linkage method calculates the average distance between all pairs of factors in two clusters, merging the clusters with the smallest average distance first. The result of the hierarchical clustering is visualized in a dendrogram (see Figure 1), a tree-like diagram that illustrates how clusters are merged at different distance thresholds. For instance, if satisfaction with benefits and satisfaction with salary are closely related, they will be merged early in the dendrogram.
@@
Additionally, we made a slight adjustment by separating the overall satisfaction factor from CF3, treating it as a distinct composite factor. This adjustment led to a final set of seven composite factors, as shown in Table 3. For example, satisfaction with benefits and salary was highly correlated, so we combined these factors into the composite factor compensation. When combining factors, we averaged their scores; for instance, if the satisfaction scores for benefits and salary were $3$ and $4$, respectively, the score for compensation would be $(3+4)/2=3.5$.
@@
Figure 1. Dendogram of clustering job satisfaction factors into groups.
Table 3. Composite Factors of Practitioners’ Job Satisfaction.
@@
5.4. Ordinal Logistic Regression Analysis
After establishing the job satisfaction clusters, we aimed to understand how fairness perception or fairness contexts are associated with the probability of job satisfaction being in a higher or lower category. To achieve this, we performed an ordinal logistic regression analysis. Ordinal logistic regression predicts the probability of an observation falling into one of several ordered categories, making it ideal for analyzing outcomes like job satisfaction levels (Agresti, 2010).
@@ 5.4. Ordinal Logistic Regression Analysis
5.5. Moderation Analysis
After performing the ordinal logistic regression analysis, we explored how the observed relationships differ across demographic groups by employing moderation analysis through RQ2 and RQ3.1. Moderation analysis tests whether a variable, such as a demographic factor, influences the direction and/or strength of the relationship between an independent variable (e.g., perception of fairness) and a dependent variable (e.g., job satisfaction) (Baron and Kenny, 1986). For example, suppose we find that perceptions of procedural fairness are positively associated with job satisfaction. A moderation analysis might examine whether this relationship is stronger for women than for men.
@@ 5.5. Moderation Analysis
In our study, we specifically examined the potential moderating effects of various demographic factors, as described in Section 6.1. Due to the imbalance in the distribution of demographic categories among our respondents (as shown in Table 4), initially, we were only able to effectively analyze the moderating effects for four demographic groups: gender, ethnicity representation, and work limitation. Subsequently, we simplified the categories for the remaining demographic groups to improve the analysis. For instance, ethnicity was grouped into Asian and Non-Asian categories, and work experience was divided into less than 3 years and more than 3 years. This refinement allowed us to effectively analyze the moderating effects for one additional demographic group: work experience.
However, the limited number of respondents in other remaining demographic categories, namely age, ethnicity, role and role nature restricted our ability to conduct reliable moderation analyses for these groups. This limitation is due to the potential for biased estimates caused by imbalanced distribution within the sample (Zhang and Wang, 2017).
6. Findings
6.1. Respondents Demographics
The demographic data is summarized in Table 4. There were 80 male respondents and 25 female respondents. The age distribution shows that 100 respondents were under 40 years old, while 7 were 40 years or older. Regarding ethnicity, the majority of respondents identified as Asian (79) and White (18). In terms of job roles, most respondents were Software Developers/Engineers (72) and Software Testers/QA Engineers (12). For work experience, 11 respondents had $\leq$ 1 year of work experience, 40 had $1-3$ years, 26 had $3-5$ years, 14 had $5-7$ years, and 16 had over 7 years of experience. Regarding work limitations, 55 respondents reported no limitations, 43 indicated some form of limitation.
Table 4. Respondents’ Demographics
@@ 6.1. Respondents Demographics
6.2. Survey Responses Summary
@@
6.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
The results of the analysis are summarized in the first section of Table 6. We considered relationships to be impactful if they were statistically significant ($p<0.05$) and had OR exceeding 50%, to ensure that the identified factors not only demonstrate a reliable statistical relationship but also have a practically meaningful impact on job satisfaction (see (fig, 2024)). Our analysis demonstrates that perceptions of fairness across various dimensions significantly shape job satisfaction among software practitioners.
@@ 6.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
Table 6. Ordinal Regression Analysis Result
@@
Takeaway 1: Fairness across all four dimensions; distributive, procedural, interpersonal, and informational, significantly enhances software practitioners’ overall job satisfaction and satisfaction with job security.
6.4. RQ2: Does the relationship between fairness perceptions and job satisfaction change based on demographic groups? If so, how?
In this RQ, we aim to determine whether a software practitioner’s demographic background influences how their perceptions of fairness affect job satisfaction using moderation analysis. The demographic groups analyzed include gender, ethnicity representation, work experience and work limitation. Similar to the previous analysis, we considered a demographic group to be influential if the effect was statistically significant ($p<0.05$) and the effect size exceeded 50% (see (fig, 2024)).
Our analysis reveals that the relationship between fairness perceptions and job satisfaction does indeed vary across demographic groups. Table 7 (first section) presents the key findings, with different emojis indicating how the perception of fairness influences job satisfaction among practitioners in the demographic groups represented by those emojis.
@@
Our analysis also unveils that distributive, procedural and informational fairness have a significant impact on female developers’ satisfaction with their manager (MNG). All distributive fairness items (D1-D4) have an average 62% stronger effect on female practitioners compared to male practitioners. Additionally, other items related to procedural fairness, such as accurate information (P3), free of bias processes (P4), consistency in applying processes (P6), and appealing decisions (P7), as well as tailored communication (Inf4) from informational fairness, also contribute to female developers’ satisfaction with their manager. This suggests that to increase female developers’ satisfaction with their managers, organizations should focus on ensuring fair outcomes, accurate information, unbiased processes, and tailored communication.
The analysis also revealead that female developers’ satisfaction with job security (SEC) is influenced by all four dimensions of fairness perceptions. Key factors include receiving outcomes that accurately reflect their work (D2) and effort (D4), the ability to voice opinions (P2), access to accurate information (P3), free of bias processes (P4), the ability to appeal decisions (P7), being treated with respect (Int2) and dignity (Int3), and receiving comprehensive explanations (Inf5). To improve job security satisfaction for female developers, organizations should ensure that performance outcomes are fair and reflective of their efforts, foster a respectful environment, and maintain clear, unbiased communication.
@@
Takeaway 2: The impact of fairness perceptions on job satisfaction is stronger for female practitioners, underrepresented practitioners, less experienced practitioners, and those with work limitations. This suggests that different demographic groups have distinct needs.
@@
6.5. RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
The results of the regression analysis are summarized in the third section of Table 6. We found positive and significant correlations between job satisfaction and fairness experiences in specific SE contexts (see (fig, 2024)).
@@ 6.5. RQ3: How do fairness perceptions influence job satisfaction in specific SE contexts?
Our analysis confirms that fairness perceptions influencing job satisfaction do indeed differ across demographic groups within SE contexts (see (fig, 2024)). Table 7 (last section) illustrates that fairness perceptions impacting job satisfaction do vary across demographic groups in SE contexts. Different emojis in the table indicate how perceptions of fairness influence job satisfaction among practitioners within the demographic groups represented by those emojis.
@@
This suggests that female developers feel more secure in their jobs when organizations ensure that performance outcomes are fair and reflective of their contributions (Lucht, 2016), foster a respectful work environment (O’farrell and Harlan, 1982), and maintain clear and unbiased communication (Iseke and Pull, 2019). These can be applied for example, when a female developer is assigned to work (i.e., allocation of work) on a critical project with tight deadlines (i.e., demand), and that workload distribution is monitored (i.e., policy) (Pater et al., 2009).
@@
We also found that unbiased processes (P4) is particularly important for underrepresented practitioners’ overall satisfaction (ALL), especially in the context of policy (C4). This suggests that ensuring unbiased processes likely helps underrepresented developers feel that they are being treated fairly, which can mitigate feelings of marginalization or exclusion (Dover et al., 2019).
@@
This suggests that how working hours are managed plays a significant role in shaping the performance satisfaction of less experienced practitioners. It could reflect how an organization or manager considers and respects a practitioner’s time and personal boundaries in relation to their workload. This includes balancing work and personal life, ensuring reasonable working hours, and handling periods of overtime with mutual respect. Less experienced practitioners may be more sensitive to these factors, as they are still establishing themselves professionally (Rastogi et al., 2015; Ando et al., 2015), making respectful management of their time essential to their perception of fairness and performance satisfaction.
@@
This could be because, at this early stage in their careers, less experienced practitioners are seeking affirmation that their contributions are valued, and they rely more heavily on external validation, such as feedback, promotions, or compensation, to gauge their progress (Smith et al., 2013; Li et al., 2011). In contexts such as policy, working hours, allocation of work, and demand, these practitioners are still learning how the workplace functions and what is expected of them. When their efforts lead to clear, positive outcomes, they are more likely to feel valued and motivated. On the other hand, if their contributions are not recognized or they feel disrespected, it may lead to dissatisfaction, as they might interpret this as a lack of growth.
@@
Takeaway 4: Regarding job satisfaction, female practitioners are more impacted by fairness in policy implementation and evaluations. Underrepresented practitioners value fairness in policy implementation and during high-demand situations. Less experienced practitioners place greater emphasis on fairness when navigating working hours, while practitioners with work limitations prioritize fairness during demanding situations.
7. Discussion
7.1. On Overall Satisfaction and Job Security
This study found that all four dimensions of fairness, namely distributive, procedural, interpersonal, and informational fairness, significantly influence overall job satisfaction and satisfaction with job security. Job security in this context relates to the perceived stability of one’s employment.
In the SE field, prior research has emphasized the importance of various factors such as autonomy, task variety, and feedback in influencing job satisfaction, particularly in agile environments (Tessem and Maurer, 2007). Autonomy, the degree to which software engineers can control how they work, and feedback on performance are particularly important in shaping their sense of satisfaction and productivity. Similarly, Storey et al. (Storey et al., 2021) highlighted appreciation, rewards, impactful work, and work culture as important factors contributing to overall job satisfaction for developers.
The closest concept to job security in the SE literature is turnover. Turnover in software development refers to the resignation or replacement of software developers, which may lead to severe problems like project postponement and failure (Ma et al., 2019). In relation to turnover, studies have found that factors such as satisfaction with supervision, workplace morale, and managerial support significantly impact turnover intentions. Research on the impact of developer turnover in open-source software projects highlights the negative impact of high developer turnover, especially among external newcomers, on software quality (Foucault et al., 2015).
Agile methods have also been linked to higher job satisfaction and lower turnover rates. A study showed that team members using agile methods reported greater job satisfaction and lower turnover intentions compared to those on non-agile teams (Melnik and Maurer, 2006). Finally, job satisfaction, onboarding success, and workplace relationships are key factors influencing turnover intentions among software professionals. Training and support for new hires are crucial for retention in these contexts (Sharma and Stol, 2020).